# OpsEval: A Comprehensive Benchmark Suite for Evaluating Large Language Models’ Capability in IT Operations Domain

### v4

@@ A.2 Dependance Filtering Keyword List
A.3 Task Types of Questions
We categorize all questions in OpsEval into 8 tasks. The details of each task are as follows:
  • General Knowledge pertains to foundational concepts and universal practices within the Ops domain.
  • Fault Analysis and Diagnostics focuses on detecting and addressing discrepancies or faults within a network or system, and deducing the primary causes behind those disruptions.
  • Network Configuration revolves around suggesting optimal configurations for network devices like routers, switches, and firewalls to ensure their efficient and secure operations.
  • Software Deployment deals with the dissemination and management of software applications throughout the network or system, verifying their correct installation.
  • Monitoring and Alerts harnesses monitoring tools to supervise network and system efficiency and implements alert mechanisms to notify administrators of emerging issues.
  • Performance Optimization is centered on refining the network and system for peak performance and recognizing potential enhancement areas.
  • Automation Scripts involves the formulation of automation scripts to facilitate processes and decrease manual intervention for administrators.
  • Miscellaneous comprises tasks that do not strictly adhere to the aforementioned classifications or involve a combination of various tasks.
A.4 Ability Levels of Questions
Different questions require different levels of ability to answer. We classify all questions in OpsEval into 3 categories. The details of each ability are as follows:
  1. 1.
    Knowledge Recall: Questions under this category primarily test a model’s capacity to recognize and recall core concepts and foundational knowledge. Such questions are akin to situations where a professional might need to identify a standard procedure or recognize a well-known issue based solely on previous knowledge.
  2. 2.
    Analytical thinking: These questions demand more than mere recall. They necessitate a deeper level of thought, expecting the model to dissect a problem, correlate diverse pieces of information, and derive a coherent conclusion. It mirrors real-world scenarios where professionals troubleshoot complex issues by connecting various dots and leveraging their comprehensive understanding.
  3. 3.
    Practical Application: These questions challenge a model’s ability to apply its foundational knowledge or analytical conclusions to provide actionable recommendations for specific scenarios. It epitomizes situations where professionals are expected to make decisions or suggest solutions based on in-depth analysis and expertise.
Figure 6: Three examples of the processed questions
Figure 6 illustrates examples in our question set, shedding light on our classification methodology.
A.5 Prompt and Formatting of Questions
Figure 7: Three examples of the processed questions
A.6 An Example of Subjective Questions
Figure 8: An example of the saved subjective questions
A saved subjective question in OpsEval is presented in Figure 8, which contains not only the raw question but also its type of task.
Figure 9: An example of building the prompt of subjective questions.
As shown in Figure 9, we combine the task and ability of each question with the question itself as the prompt for LLMs.
A.7 Automated QA generation
During the data collection process, we have experimented automating question-answer generation. We first sampled the QA pairs and manually assessed their accuracy and domain relevance. Later, we used typical manual evaluation examples for few-shot learning, enabling GPT to evaluate QA pairs based on our evaluation criteria automatically. Directly generated question-answers tend to be simple judgment or concept questions rather than reasoning questions that better demonstrate the model’s capabilities and knowledge density. Our goal is to ensure that while the topics of the questions remain relevant to the seed questions, their specific content is distinct from the original questions. By maintaining the overarching framework in the Ops domain, we can expand the number and types of questions, enabling a more comprehensive evaluation of model capabilities. Additionally, we can incorporate external knowledge during the data generation, continually enhancing our ability to evaluate new content.
Appendix B Additional details of experiments
B.1 Detailed Information of LLMs Evaluated
Table 9: Models evaluated in this paper. The “access” column in the table shows whether we have full access to the model weights or can only access them through API.
Model Creator #Parameters Access License
GPT-4/GPT-3.5-turbo OpenAI undisclosed API Proprietary
ERNIE-Bot-4.0 Baidu undisclosed API Proprietary
GLM4/GLM3-turbo Tsinghua Zhipu undisclosed API Proprietary
Meta-LLaMA-3 Meta 8B Weights Llama 3 Community
LLaMA-2 Meta 7/13/70B Weights Llama 2 Community
Qwen-Chat Alibaba Cloud 7/14/72B Weights Qianwen LICENSE
Qwen1.5-Chat Alibaba Cloud 14B Weights Qianwen LICENSE
InternLM2-Chat Shanghai AI Laboratory 7/20B Weights Apache-2.0
DevOps-Model-Chat CodeFuse 14B Weights Apache-2.0
Baichuan2-Chat Baichuan Intelligence 13B Weights Apache-2.0
ChatGLM3 Tsinghua Zhipu 6B Weights Apache-2.0
Mistral Mistral 7B Weights Apache-2.0
Gemma Google 2/7B Weights Gemma license
GPT-4 (OpenAI, 2023a) is a large multimodal model (accepting image and text inputs, emitting text outputs) that, while less capable than humans in many real-world scenarios, exhibits human-level performance on various professional and academic benchmarks. It is recognized as the strongest lanuage model currently. ChatGPT (OpenAI, 2022) is an earlier AI-powered language model developed by OpenAI which is built upon GPT-3.5. We use the GPT-3.5-turbo version in our experiments. LLaMA 2 (Touvron, et.al., 2023) is a second-generation open-source LLM from Meta which is very popular due to its open-source feature. It has the ability to process multiple languages including Chinese. We evaluate three weights (70B, 13B and 7B as shown in 9) of LLaMA 2.
Although LLaMA 2 is able to process Chinese input, it has a small Chinese vocabulary so that its abitilty of understanding and generating Chinese text is limited. As a result, we evaluate some Chinese-oriented LLMs which are published by institutions in China. ERNIE-Bot 4.0 (202, 2024) is the latest self-developed language model released by Baidu. As claimed by Baidu, ERNIE-Bot 4.0 rivals OpenAI’s GPT-4. Qwen (202, 2023) (abbr. Tongyi Qianwen) is a series of LLMs developed by Alibaba Cloud. And Qwen-Chat is a series of large-model-based AI assistant trained with alignment techniques based on the pretrained Qwen. We evaluate three weights (72B, 14B and 7B as shown in 9) of Qwen-Chat. Baichuan2-13B-Chat (Baichuan, 2023) is aligned chat model based on Baichuan2-13B-Base (Baichuan, 2023) which is an open-source LLM published by Baichuan Intelligence. GLM (Du et al., 2022), developed by Tsinghua Knowledge Engineering Group, is a General Language Model pretrained with an autoregressive blank-filling objective and can be finetuned on various natural language understanding and generation tasks. Based on GLM, Zhipu AI released GLM4 (the newest version of GLM model) (Zeng et al., 2022) and GLM3 (the third version of GLM model). For GLM3, we use GLM3-turbo (Zeng et al., 2022) version and ChatGLM3-6B (Zeng et al., 2022) in our experiments. InternLM2-Chat-20B and InternLM2-Chat-7B (InternLM_Team, 2023), recently developed by Shanghai AI Laboratory, are multi-lingual models based on billions of parameters through multi-stage progressive training on over trillions of tokens. Furthermore, we evaluate DevOps-Model-14B-Chat (AI, 2024a), an open source Chinese DevOps oriented models, mainly dedicated to exerting practical value in the field of DevOps.Gemma  (Gemma_Team et al., 2024) is a family of lightweight, state-of-the-art open models based on Gemini technology from Google DeepMind. Trained on up to 6T tokens, Gemma achieves excellent language understanding and reasoning capabilities. We conducted an evaluation of Gemma-2b and Gemma-7b to investigate the effectiveness of Gemma with different weights.
In general, since some models (among them GPT-4, GPT-3.5-turbo, ERNIE-Bot-4.0, GLM4, GLM3-turbo) are not locally available, we evaluate them via API calls. For the remaining models, we perform local inference during evaluation.
Table 10: GPTQ models for LLaMA-2-70B
Model Size #GPTQ Dataset Disc
LLaMA-2-70B 140GB / Raw LLaMA-2-70B model.
LLaMA-2-70B-Int4 35.33GB wikitext 4-bit quantization model.
LLaMA-2-70B-Int3 26.78GB wikitext 3-bit quantization model.
B.2 Prompts
Figure 10: An example of zero-shot evaluation in the CoT setting.Black font represents prompts in English. Purple font represents prompts in Chinese. Red font represents the model’s output in Chinese. Dark red font represents the model’s output in English.
Figure 11: An example of few-shot evaluation in the CoT setting.Black font represents prompts in English. Purple font represents prompts in Chinese. Red font represents the model’s output in Chinese. Dark red font represents the model’s output in English.
For zero-shot evaluation in the CoT setting, we get the answer of LLMs in two rounds. Firstly, by adding a ’Let’s think step by step.’ after the question, LLMs will output its reasoning result. Secondly, we compose the final prompt of the question and the reasoning result in whole as the input of LLMs to get the final answer. An example is shown in Figure 10. For few-shot evaluation in the CoT setting, We make an analysis of each option of the question as a reasoning process, and craft three Q-A examples with CoT reasoning process in answers. An example is shown in Figure 11.
B.3 Compute and Resources Used for Experiments
During our OpEval experiments evaluating different LLMs, we utilize an 8 Nvidia A800-80GB GPU cluster to run inference on models with available weights. For models with API access, we perform inference using CPUs.
B.4 Overview Performance on Different Test Sets
Table 11: LLMs’ overall performance on wired network operations test set
Model English Test Set Chinese Test Set
Zero-shot 3-shot Zero-shot 3-shot
Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC
GPT-4 / / / / / / 88.70 / / / / / / / 86.00 /
Qwen-72B-Chat 70.41 70.50 72.38 72.56 70.32 70.32 70.13 70.22 65.77 65.86 68.13 68.30 69.40 69.40 69.99 70.08
GPT-3.5-turbo 66.60 66.80 69.60 72.00 68.30 68.30 70.90 72.50 58.40 58.60 64.80 67.60 59.20 59.70 65.20 67.40
ERNIE-Bot-4.0 61.15 61.15 70.00 70.00 60.00 60.00 70.00 70.00 67.54 67.54 71.96 71.96 72.00 72.00 78.00 78.00
Qwen1.5-14B-Chat 54.90 34.88 64.09 60.82 52.23 65.55 59.54 47.08 54.04 45.18 62.56 59.12 58.78 61.10 63.43 52.5
Devops-Model-14B-Chat 30.69 30.59 55.77 63.63 63.85 61.96 41.15 44.01 47.59 46.57 52.52 56.01 62.07 60.08 50.59 55.79
Qwen-14B-Chat 43.78 47.81 56.58 59.40 62.09 59.70 49.06 55.88 48.35 48.81 55.35 57.40 58.53 56.12 52.12 54.99
LLaMA-2-13B 41.80 46.50 53.10 58.70 53.30 53.00 56.80 61.00 29.70 31.60 51.60 57.00 39.60 38.90 48.00 50.60
Gemma-7B 25.09 25.09 50.86 50.86 59.12 59.12 50.77 50.77 31.58 31.58 47.59 47.59 34.68 34.68 48.88 48.88
LLaMA-2-70B-Chat 25.29 25.29 57.97 58.06 52.97 52.97 58.55 58.55 38.55 38.55 57.49 57.49 49.09 49.09 48.57 48.57
Internlm2-Chat-20B 56.36 56.36 26.18 26.18 60.48 60.48 45.10 45.10 57.49 57.49 57.14 57.14 59.12 59.12 50.77 50.77
Internlm2-Chat-7B 49.74 49.74 56.19 56.19 48.20 48.20 49.74 49.74 57.49 57.49 57.14 57.14 59.12 59.12 50.77 50.77
LLaMA-2-7B 39.50 40.00 45.40 49.50 48.20 46.80 52.00 55.20 29.80 30.20 50.10 55.60 38.60 40.80 45.60 50.40
Qwen-7B-Chat 45.90 46.00 47.30 50.10 52.10 51.00 48.30 49.80 29.60 29.90 50.60 53.50 50.40 46.90 46.90 47.70
Baichuan2-13B-Chat 37.90 38.30 42.70 46.60 51.90 51.60 44.50 47.45 44.60 45.40 41.60 44.30 45.60 45.70 43.90 46.70
Note: The best accuracy of each language for each LLM is in bold font.
Table 12: LLMs’ overall performance on 5G communication operations test set
Model English Test Set Chinese Test Set
Zero-shot 3-shot Zero-shot 3-shot
Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC
GPT-4 / / 56.30 65.49 / / 59.62 63.54 / / 57.19 62.11 / / 61.55 65.68
Qwen-72B-Chat 53.19 53.19 55.25 55.52 58.13 58.13 58.72 58.99 64.79 64.79 65.79 65.72 70.19 70.19 68.31 68.38
InternLM2-Chat-20B 39.10 39.10 37.70 37.70 47.70 47.70 33.50 33.50 44.60 44.60 47.00 47.00 62.20 62.20 38.30 38.30
Qwen-14B-Chat 33.71 36.25 41.24 42.51 51.19 50.39 57.18 59.18 41.71 41.44 45.58 47.98 53.52 49.92 54.72 58.85
DevOps-Model-14B-Chat 31.04 30.51 42.84 47.37 52.25 49.38 45.90 47.23 41.04 42.70 48.71 53.57 56.85 57.25 51.30 54.29
ERNIE-Bot-4.0 43.66 43.66 51.99 51.99 44.00 44.00 50.00 50.00 45.99 45.99 48.98 48.98 46.00 46.00 54.00 54.00
LLaMA-2-70B 23.64 23.64 39.31 39.31 38.98 39.12 47.90 47.90 24.38 24.38 43.63 43.63 44.65 44.65 48.84 48.84
Mistral-7B 26.91 26.91 30.65 30.65 40.52 40.52 46.84 46.84 1.27 1.27 42.05 42.05 30.72 30.72 46.44 46.44
InternLM2-Chat-7B 36.80 36.80 31.70 31.70 46.30 46.30 36.90 36.90 38.80 38.80 44.60 44.60 46.00 46.00 35.80 35.80
Gemma-7B 23.10 23.10 34.40 34.40 21.40 21.40 33.10 33.10 27.30 27.30 35.40 35.40 17.30 17.30 44.50 44.50
LLaMA-2-13B 15.62 18.32 29.88 34.45 23.16 29.14 37.59 44.3 25.43 27.16 29.17 29.99 36.56 36.15 37.70 39.02
GPT-3.5-turbo 34.92 34.82 38.53 43.50 39.40 39.19 40.93 42.58 36.98 36.83 37.95 39.25 39.17 39.77 41.93 42.15
Qwen-7B-Chat 33.85 33.74 32.45 34.10 32.91 32.70 36.65 36.65 36.27 36.50 33.27 33.51 42.22 40.59 31.28 31.46
ChatGLM3-6B 30.40 30.40 30.70 30.70 26.90 26.90 37.20 37.20 32.60 32.60 35.40 35.40 28.30 28.30 40.90 40.90
Baichuan2-13B-Chat 14.10 15.30 24.10 25.80 32.30 33.10 25.60 27.70 35.64 35.91 30.59 30.52 34.65 35.6 30.21 32.05
LLaMA-2-7B 19.14 21.62 25.70 27.11 21.38 24.85 32.38 34.83 23.57 23.47 27.65 29.26 30.30 30.03 30.98 31.93
Gemma-2B 20.10 20.10 24.20 24.20 31.20 31.20 35.50 35.50 25.60 25.60 28.30 28.30 19.10 19.10 35.50 35.50
Note: The best accuracy of each language for each LLM is in bold font.
Table 13: LLMs’ overall performance on database operations test set
Model English Test Set Chinese Test Set
Zero-shot 3-shot Zero-shot 3-shot
Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC
GPT-4 / / 59.02 64.56 / / 58.35 62.58 / / 59.38 65.17 / / 44.06 48.09
InternLM2-Chat-20B / / 59.21 59.21 / / / / / / / / / / / /
ERNIE-Bot-4.0 43.80 43.80 47.14 47.14 46.00 46.00 54.0 54.0 48.56 48.56 50.64 50.64 48.00 48.00 54.0 54.0
Gemma-7B 14.29 14.29 30.99 30.99 2.60 2.60 43.86 43.86 19.32 19.32 53.95 53.95 18.51 18.51 5.20 5.20
Qwen-72B-Chat 47.28 47.48 48.09 48.09 49.70 49.70 43.46 43.66 48.29 48.49 49.50 49.70 49.70 49.70 45.27 44.87
GPT-3.5-turbo 38.63 38.83 40.04 42.05 36.62 37.63 42.66 43.86 36.42 35.81 39.24 43.26 39.84 39.44 27.16 27.77
Qwen-14B-Chat 24.95 28.37 33.00 36.62 27.97 28.37 27.97 24.14 27.57 27.57 32.39 36.02 40.04 35.41 30.38 33.40
DevOps-Model-14B-Chat 25.15 26.96 35.41 38.83 33.20 34.81 27.36 27.36 24.75 22.74 28.37 27.77 36.62 37.02 27.57 26.36
LLaMA-2-70B 19.72 19.72 27.97 27.97 26.56 26.56 32.6 32.6 15.29 15.29 34.81 34.81 26.76 26.76 33.80 33.80
Qwen-7B-Chat 18.91 19.11 22.13 23.94 26.76 25.55 34.81 34.81 18.51 17.71 27.36 28.37 29.78 29.58 33.60 33.60
LLaMA-2-13B 16.10 20.32 23.94 29.58 20.12 22.33 24.35 33.80 23.94 24.35 29.58 31.99 24.55 26.76 21.13 20.72
LLaMA-2-7B 22.13 23.74 23.74 26.56 19.32 20.52 28.77 33.60 20.72 20.72 27.16 27.97 21.53 18.51 18.31 17.91
Mistral-7B 17.10 17.10 26.76 26.76 31.19 31.19 27.97 27.97 0.20 0.20 26.76 26.76 10.26 10.26 32.19 32.19
InternLM2-Chat-7B 27.16 27.16 28.17 28.17 29.98 29.98 30.18 30.18 28.57 28.57 31.79 31.79 30.78 30.78 31.19 31.19
ChatGLM3-6B 20.93 20.93 25.15 25.15 24.75 24.75 29.18 29.18 21.33 21.33 28.97 28.97 21.73 21.73 29.58 29.58
Baichuan2-13B-Chat 17.10 19.11 18.71 22.94 25.96 26.56 20.93 24.55 25.75 25.55 20.12 21.33 27.77 26.76 22.74 24.75
Gemma-2B 16.90 16.90 19.52 19.52 16.10 16.10 24.75 24.75 18.51 18.51 24.95 24.95 21.53 21.53 27.77 27.77
Note: The best accuracy of each language for each LLM is in bold font.
In Table 11, Table 12 and Table 13, we present overview performance of different LLMs on the 3 test sets in OpsEval, including Wired Network Operations, 5G Communication Technology Operations and Database Operations.
B.5 Performance on Different Quantization Models
Figure 12: LLaMA-2-70B’s performance of different quantization parameters. Both zero-shot and few-shot evaluations have been conducted on Wired Network Operations test set under the naive setting.
Figure 12 shows the accuracy of LLaMA-2-70B of different quantization parameters on objective questions, English and Chinese questions respectively. We do both zero-shot and few-shot evaluation with the naive setting.
[…]

### v5

@@ 2.2. Preprocessing
Question Categorization: We devise a categorization that captures many tasks that professionals confront in practical applications. The categorization process consists of two steps: automated screening and manual review. We first use GPT-4 for topic modeling to gain rough insights about the dataset and determine the relevance of each question to Ops, which resulted in more than 20 tasks but had an imbalanced distribution. We then involved dozens of experts during the manual review process to categorize the questions into eight tasks and three abilities.
Tasks. The details of each task are as follows. General Knowledge pertains to foundational concepts and universal practices within the Ops domain. Fault Analysis and Diagnostics focuses on detecting and addressing discrepancies or faults within a network or system, and deducing the primary causes behind those disruptions. Network Configuration revolves around suggesting optimal configurations for network devices like routers, switches, and firewalls to ensure their efficient and secure operations. Software Deployment deals with the deployment and management of software applications throughout the network or system. Monitoring and Alerts harnesses monitoring tools to supervise network and system efficiency and implements alert mechanisms to notify administrators of emerging issues. Performance Optimization is centered on refining the network and system for peak performance and recognizing potential enhancement areas. Automation Scripts involves the formulation of automation scripts to facilitate processes and decrease manual intervention for administrators. Miscellaneous comprises tasks that do not strictly adhere to the aforementioned classifications or involve a combination of various tasks.
Abilities. Different questions require different levels of ability to answer. Knowledge Recall primarily test a model’s capacity to recognize and recall core concepts and foundational knowledge, which are akin to situations where professionals identifies a standard procedure or recognize an issue based solely on previous knowledge. Analytical thinking necessitates a deeper level of thought, expecting the model to dissect a problem, correlate diverse pieces of information, and derive a coherent conclusion. It mirrors scenarios where professionals troubleshoot complex issues by leveraging their comprehensive understanding. Practical Application challenges a model to apply its knowledge or analytical conclusions to provide actionable recommendations for specific scenarios. It mirrors situations where professionals make decisions or suggest solutions based on in-depth analysis and expertise.
Figure 2. Three examples of the processed questions.
[…]
[Three examples of the processed questions.]Three examples of the processed questions.
Figure 2 illustrates examples in our question set, demonstrating our classification methodology. The distribution of the questions across the tasks and ability levels is shown in Table 1(b).
Manual Review: In the manual review step, we asked Ops experts from the industry to inspect the results of the previous three automated steps, including confirming duplicate and invalid questions and examining the classification results of GPT-4. In our work, an expert is defined as an individual with ten or more years of professional experience in their field, whether as an employee or a researcher. Experts were also asked to drop the questions unrelated to the Ops field. We split the dataset by n-folds and ensure each fold has at least two experts to review. The review process followed a standardized annotation guideline, which is available in our dataset repository. As listed in Table 1(a), this quality enhancement process resulted in a refined test set of approximately 7,000 multi-choice and 2,000 question-answering questions.
Figure 3. The FAE-Score pipeline.
[…]
[The FAE-Score pipeline.]The FAE-Score pipeline.
2.3. Evaluation Settings
Table 3. Models evaluated in this paper.
Model #Parameters Access License
GPT-4/3.5-turbo undisclosed API Proprietary
ERNIE-Bot-4.0 undisclosed API Proprietary
GLM4/GLM3-turbo undisclosed API Proprietary
Meta-LLaMA-3 8B Weights Llama 3 Community
LLaMA-2 7/13/70B Weights Llama 2 Community
Qwen-Chat 7/14/72B Weights Qianwen LICENSE
Qwen1.5-Chat 14B Weights Qianwen LICENSE
InternLM2-Chat 7/20B Weights Apache-2.0
DevOps-Model 14B Weights Apache-2.0
Baichuan2-Chat 13B Weights Apache-2.0
ChatGLM3 6B Weights Apache-2.0
Mistral 7B Weights Apache-2.0
Gemma 2/7B Weights Gemma license
Claude-3-Opus undisclosed API Proprietary
Qwen2-Instruct 7/72B Weights Qianwen LICENSE
Question-answering questions are evaluated using a metric designed specifically for OpsEval, called FAE-Score, which is explained in detail in the subsequent section. Additionally, we perform expert evaluations and calculate BLEU (Papineni et al., 2002), ROUGE (Lin, 2004) and RAGAS (Es et al., 2024) scores for comparison purposes, as reference to validate the accuracy of FAE-Score.
We use the same three criteria to evaluate the responses of various models for both FAE-Score and Expert Evaluation:
In Expert Evaluation, we asked experts to score it between 0 and 3 for each criterion. During the scoring, the raw question, the detailed answer and its key points, and the output of an anonymous model are given at each iteration.
Prompting Techniques. We use various settings to evaluate LLMs on OpsEval to get a comprehensive overview of their performance. We evaluate LLMs in zero and few-shot (3-shot) settings. For each setting, we evaluate LLMs in four sub-settings of prompt engineering, that is, naive answers (Naive), self-consistency (SC) (Wang et al., 2023c), chain-of-thought (CoT) (Wei et al., 2023), self-consistency with chain-of-thought (CoT+SC). We set the number of queries in SC to 5.
Models. We evaluate popular LLMs covering different weights from different organizations. The model selection was guided by specific criteria: We aimed to include the latest and most advanced large language models, with a particular focus on those capable of handling Chinese input. The detailed information of all 24 LLMs can be found in Table 3.
2.4. FAE-Score
Figure 3 shows the basic pipeline of our designed QA metric, FAE-Score. Here, we elaborate each evaluation methodology of each criterion.
Fluency. In Ops settings, the fluency of a model’s output is crucial because the results are intended for human consumption by technical personnel. Unlike other generic benchmarks, the tasks in the Ops domain require clear and unambiguous communication, as the model’s outputs may guide decision-making in production scenarios. To evaluate fluency in model outputs, we adapted the scoring rubrics methodology mentioned in Kim et al. (2024). We use Qwen2-72B-Instruct as the evaluation model, for its strong performance in general language generation (QwenLM, 2023) and its consistent multilingual capabilities. We assess the fluency of various model outputs based on grammar, coherence, clarity, appropriateness of style, and answer completeness, as shown in Figure 4.
Figure 4. Scoring rubrics for Fluency metric.
[…]
[Scoring rubrics for Fluency metric.]Scoring rubrics for Fluency metric.
Accuracy. Traditional metrics such as BLEU and ROUGE fall short in this vertical domain because they often fail to capture the key factual content within long-form responses. This results in inflated scores due to irrelevant word matches, making these metrics insufficient for accuracy evaluation in the highly specialized and knowledge-driven Ops context. To address these shortcomings, we take inspiration from Es et al. (2024), using a keyword extraction method to evaluate the accuracy of model outputs. A judge model (OpenAI, 2023a) is then employed to match the keywords from the model’s response with the keywords from the standard answer. The final accuracy score is calculated by determining the F1-Score, which balances precision and recall for the matched keywords.
(1) $\textbf{Accuracy}=2\cdot\frac{\text{P}\cdot\text{R}}{\text{P}+\text{R}}$
(2) P=#Matched Keywords#Keywords in Model $\text{P}=\frac{\#{\text{Matched Keywords}}}{\#{\text{Keywords in Model Output}}}$
(3) R=#Matched Keywords#Keywords in Ground $\text{R}=\frac{\#{\text{Matched Keywords}}}{\#{\text{Keywords in Ground Truth}}}$
Figure 5. LLMs’ overall performance on Wired Network Operations English test set (3-shot). Models are ranked based on their mean accuracy among different settings. The error bars represent the variance in the model’s accuracy across different prompting techniques.
[…]
[LLMs’ overall performance on Wired Network Operations English test set (3-shot).]LLMs’ overall performance on Wired Network Operations English test set (3-shot). Models are ranked based on their mean accuracy among different settings. The error bars represent the variance in the model’s accuracy across different prompting techniques.
Evidence. Model responses must not only be accurate but also well-supported by relevant, authoritative information. To evaluate the evidence behind a model’s response, we implement a ROUGE-based method to measure the overlap between the generated output and the content of related documents retrieved through similarity search. We used bge-large-zh (Xiao et al., 2023) for document embedding and FAISS (Douze et al., 2024) for similarity search. By retrieving documents that closely match the question, we can assess whether the model’s response appropriately references or aligns with this external information. We use ROUGE, as a recall-oriented metric, captures how much of the content in the relevant documents is reflected in the model’s output. This ensures that the model does not simply generate plausible-sounding answers but grounds its responses in factual evidence from trusted sources.
(4) $\textbf{Evidence}=\text{ROUGE}_{\text{Recall}}(R,D)=\frac{\#\text{Overlapping Words}}{\#\text{Words in }D}$
2.5. Open-source Policy
We released 20% of the OpsEval dataset to support research and community contributions. The subset was proportionally sampled across sources and sub-domains, with proprietary content reviewed to remove sensitive information. This sample dataset offers a clear view of question types and topics in the benchmark, helping researchers grasp the evaluation scope. It also supports local model evaluation for quicker iteration and can seed automatic QA generation (Wang et al., 2023b), enriching Ops-specific data for future development. While this subset is available for users’ self-evaluation, the complete dataset remains undisclosed. By ensuring that the test set answers are not leaked, we guarantee the reliability and non-leakage of the OpsEval benchmark.
3. Result Analysis
3.1. Overall Performance
The results of the few-shot evaluation with four settings on the Wired Network Operation test set are shown in Figure 5.  [^2] For results of the other sub-domains and settings, please check our official leaderboard website https://opseval.cstcloud.cn/content/leaderboard. While closed source models like GPT-4 and Claude-3-Opus performs well on the OpsEval benchmark, open-sourced LLMs yield generally worse evaluation results than those in general domains like MMLU (Hendrycks et al., 2021) and CEval (Huang et al., 2023). This comparison highlights the necessity of explicitly fine-tuning OpsLLM for the Ops field. Recent open-sourced models like Qwen2-72B-Chat, exhibit competitive performance in multi-choice questions, thanks to their fine-tuning process and the quality of their training data. Furthermore, we observed significant variability in how different LLMs respond to various prompt engineering techniques. Given the critical importance of stability in the Ops domain, it is essential to consider a model’s sensitivity to prompts when selecting foundation model. Further research into prompt engineering is needed to improve model performance and reliability in this domain.
Observations: 1) Few-shot and CoT can significantly increase performance if the model is tuned to adapt to these techniques, while SC may have little influence on highly consistent LLMs. 2) Smaller models with weaker abilities are less stable with advanced prompts. Simpler prompts work better for them.
3.2. Performance on Different Tasks and Abilities
Figure 6. LLMs’ performance on eight Ops sub-domains, eight tasks and three abilities. Each colored area presents the lower and upper bound of the corresponding parameter-size group.
[…]
[LLMs’ performance on eight Ops sub-domains, eight tasks and three abilities.]LLMs’ performance on eight Ops sub-domains, eight tasks and three abilities. Each colored area presents the lower and upper bound of the corresponding parameter-size group.
Figure 7. Heatmap of failure case distribution regarding tasks and abilities. The values represent the proportion of failure cases across all LLMs; yellower areas indicate higher failure rates.
[…]
[Failure case distribution heatmap]Heatmap of failure case distribution regarding tasks and abilities. The values represent the proportion of failure cases across all LLMs; yellower areas indicate higher failure rates.
To investigate how LLMs perform in each Ops sub-domain and each task, and to what extent they possess the general abilities, we summarize the result of different parameter-size groups of LLM in Figure 6. Regarding the eight tasks we tested, LLMs yield higher accuracy in General Knowledge tasks, while their performance drops and varies drastically in highly specialized tasks like Automation Scripts and Network Configuration, reflecting the impact of specialized corpus and domain knowledge on the performance of LLMs. By grouping LLMs by their parameter size, we find that while LLMs with 10B-30B parameters have higher accuracy in their best cases compared with LLMs with no more than 10B parameters, different 10B-20B LLMs’ performance varies drastically. To provide systematic practical lessons for researchers in the operations domain on pre-training and fine-tuning OpsLLM, we have analyzed the error rates of LLMs across the 8 tasks and 3 abilities in Figure 7. By examining the focus areas across different categories, we have identified key research targets for capability training.
Insights: 1) Among different sub-domains of Ops, 5G communication and database demand further pretraining and fine-tuning. 2) To be capable of an Ops agent, the foundation model must be able to make a connection between specialized domain knowledge.
3.3. Performance on Question-Answering
Table 4. LLMs’ performance on English network operations question-answering problems.
Model ROUGE(%) BLEU(%) RAGAS(0-10) Fluency Accuracy Evidence FAE-Total
FAE Expert FAE Expert FAE Expert FAE Expert
GPT-3.5-turbo 12.26 6.78 9.23 9.38 9.12 8.06 9.65 6.21 8.11 23.65 26.88
LLaMA-2-70B 7.74 4.2 6.04 8.69 8.25 7.71 8.79 9.08 8.98 25.48 26.02
LLaMA-2-13B 4.98 3.43 8.23 8.47 9.84 7.32 9.34 8.81 7.27 24.60 26.44
Chinese-Alpaca-2-13B 3.25 1.85 5.32 5.53 8.05 6.99 7.95 6.23 6.23 18.75 22.24
Baichuan-13B-Chat 4.76 0.35 7.93 7.16 7.98 8.71 7.84 6.66 7.31 22.53 23.13
Qwen-7B-Chat 11.82 4.33 4.92 7.63 5.82 6.42 7.27 6.57 5.37 20.62 18.47
ChatGLM2-6B 9.71 5.07 5.32 5.12 7.96 6.41 6.39 6.14 4.32 17.67 18.67
InternLM-7B 13.27 0.54 6.21 4.99 5.16 5.00 4.90 4.75 4.28 14.74 15.77
Chinese-LLaMA-2-13B 9.19 0.24 7.34 6.98 4.64 5.29 6.32 4.63 8.34 16.90 17.88
Figure 8. Case analysis on QA metrics.
[…]
[Case analysis on QA metrics.]Case analysis on QA metrics.
Table 4 presents the evaluation results of 200 question-answering English questions across four metrics: ROUGE, BLEU, RAGAS, FAE-Score, and Expert-Evaluation. To gain more insight into how different metrics perform in QA evaluation, we use Figure 8 as a case analysis. While BLEU and ROUGE are efficient in natural language comparison, they lack semantic information to determine which part of the context is more important than others. Knowing that a given benchmark evaluates QA based on BLEU/ROUGE, there is an obvious way to trick the metric: repeat patterns occurring in the question, gaining a higher possibility to match some patterns in the reference answer. Due to their lack of semantic information related to Ops and the potential hack, traditional metrics like BLEU are unsuitable for specialized benchmarks. Instead, with specialized prompting and seperately designed methodology for each criterion (Fluency, Accuracy and Evidence), FAE-Score can comprehensively evaluate models’ QA performance, with the Accuracy metric picking up those important keywords and not be influenced by repeated words that contain no useful information, and the Evidence metric checking the recall of relevant supporting contents. In Section 4, we discuss the alignment between different metrics and expert evaluation, validating the effectiveness of FAE-Score in automated QA evaluation within the Ops domain.
Insight: In specialized domains, Ops specifically, traditional NLP metrics like BLEU and ROUGE cannot comprehend the key components in the reference answer, resulting their evaluation lacking practical significance. FAE-Score is suitable for large-scale qualitative evaluations in the Ops field.
3.4. Performance on Different Quantization parameters
Figure 9. LLaMA-2-70B’s performance of different quantization parameters. Both zero-shot and few-shot evaluations have been conducted on Wired Network Operations test set under the naive setting.
[…]
[LLaMA-2-70B’s performance of different quantization parameters.]LLaMA-2-70B’s performance of different quantization parameters. Both zero-shot and few-shot evaluations have been conducted on Wired Network Operations test set under the naive setting.
We conducted experiments on different quantized versions of LLaMA-2-70B and obtained various results and conclusions. Figure 9 shows the accuracy of LLaMA-2-70B of different quantization parameters on English and Chinese questions. We do both zero-shot and few-shot evaluation with the naive setting.
[…]